• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Holmes

    Log entry time 08:02:43 on June20,2004

    Entry number 126399

    keyword=Owl Shift Summary

    00:21 R2298 ended
    00:21 R2299 production 30 uA
    01:17 R2300 production 30 uA
    02:13 R2301 production 29.5 uA
    03:11 R2302 production 30 uA
    Error messages from DAQ, problems with EPICS variables
    Target fan speed plot is strange
    04:05 R2303 production 30 uA
    Continued EPICS messages
    04:30 R2304 production 30 uA messages stop
    05:26 R2305 production 30 uA
    06:22 R2306 production 29.5 uA ended early, DAQ errors
    07:10 R2307 production 29.5 uA
    07:20 Left D1 tripped. RC paged.
    07:30 R2308 production (left D1 down)
    07:55 RC came in, magnet turns on now. Ramping up.
    08:00 D1 tripped again while ramping (around 300+ amps)